OpenStudy (anonymous):

Write an equation to translate the following, then solve it to find the number: -4 times a number increased by 16 is equal to 116

OpenStudy (anonymous):

hey @Thatgirl9513 , -4 x what number? Do we know it?

OpenStudy (anonymous):

@Thatgirl9513

OpenStudy (anonymous):

ok , just work with me here so that we quickly get to the bottom of this.

OpenStudy (anonymous):

anyway let the unknown number be\[x\]

OpenStudy (anonymous):

so \[-4 \times x + 16 =116\] \[-4x +16-16 =116-16\] \[-4x =100\] \[\frac{ -4 }{ -4}x = \frac{ 100 }{ -4 }\] \[x=-25\]

OpenStudy (anonymous):

hope u will understand

OpenStudy (anonymous):